triangle ABC is similiar to triangle DEF. Suppose a=3,b=6,c=7 ,and d=4 find e and f.
a/d= b/e = c/f
3/4= 6/e?= 7/f? why does e=8? full break down

3/4=6/8=7/f f=9 1/3 why does f= 9 1/3 full break down


3f=28
f= 9 1/3

Found 2 solutions by solver91311, RAY100:
Answer by solver91311(12126)



Already did this one. Question 191525


John


Answer by RAY100(1637)

similar triangles means that the sides are proportional.

sides a, b, c, correspond to sides d, e, f
therefore a/d = b/e = c/f or all equal scale factor between triangles.
with d=4, we can see that a / d equals 3 / 4 (given)
all other sides have same scale factor (proportion)
math is a / d = 3/4 = 6 / e = 7 / f
taken one at a time, 3 / 4 = 6 / e , cross multiplying, 3 * e = 4 * 6 = 24, or e = 24/3 = 8
likewise 3/4 = 7 / f, 3*f = 4* 7 = 28, f = 28/3 = 9 1/3